OpenStudy (anonymous):

2 5/2 - 2 3/2 = 2 i dont remember how to solve so im just guessing if i am wrong say why

OpenStudy (anonymous):

Your correct , When You have the same denominator you can add and subtract.

OpenStudy (anonymous):

2 5/2 - 2 3/2 = 9/2 - 7/2 = 2/2 = 1

OpenStudy (anonymous):

u disagree gregorv?

OpenStudy (cggurumanjunath):